OpenStudy (jkbo):

Please help Solve the system. y-10=3x 2y=6x+20 A.(2, 13) B.infinite solutions C.(10, 40) D.no solution

OpenStudy (anonymous):

first one can be written as \(y=3x+10\) substitute in to the second and solve \[2(3x+10)=6x+20\]

OpenStudy (anonymous):

since when you distribute on the left you get \[6x+20=6x+20\] you see you have an identity, true for all \(x\) so infinitely many solutions

OpenStudy (jkbo):

thank you very much @satellite73

OpenStudy (jkbo):

@satellite73 can you help me with this one also? Solve the following system by graphing and identify the point of intersection 5x-y=-9 x-y=-1 A.(–1, –2) B.(–2, –1) C.(1, 2) D.(2, 1)